Circular No. 7685

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q1 (NEAT) K. J. Lawrence, E. F. Helin, and S. Pravdo, Jet Propulsion Laboratory,

report the discovery by the Near Earth Asteroid Tracking program of a new comet on CCD images obtained with the 1.2-m Oschin Schmidt telescope at Palomar (discovery observation below).

2001 UT  α_{2000}  δ_{2000}  m_{1}
Aug. 17.20219  20^{h}57^{m}12.^{s}27  −8^{o}04′17′′.1  18.8

Lawrence notes that the object is diffuse with a nuclear condensation of diameter ∼ 3′′. Following posting on the NEO Confirmation Page, P. Pravec and P. Kušnirak (Ondřejov 0.65-m reflector) confirmed its cometary appearance on CCD images obtained on Aug. 18.9 UT, and M. Kočer (Kleť 0.57-m reflector) reports that the object is diffuse and at m_{1} = 18.0 on Aug. 18.9. T. B. Spahr, Minor Planet Center, has also identified the object in data obtained by LONEOS on July 16.2. Full astrometry and the following parabolic orbital elements appear on MPEC 2001-Q18:

T = 2001 Sept.28.947 TT  ω = 176.^{o}275
Ω = 139.109 }2000.0
q = 5.88368 AU  i = 67.900
2001 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Aug. 9 21^{h}01.^{m}01 − 6^{o}50.^{′}4 4.893 5.895 170.^{o}3 1.^{o}7 18.2
19 20 56.39 − 8 20.6 4.906 5.891 165.3 2.5 18.2
29 20 52.12 − 9 51.2 4.952 5.888 155.9 4.0 18.2
Sept. 8 20 48.46 −11 19.7 5.027 5.886 145.5 5.6 18.2
18 20 45.64 −12 43.7 5.130 5.884 135.0 6.9 18.2

SUPERNOVA 2001dm IN NGC 749 D. Baade, European Southern Observatory (ESO); C. Fransson, Stock-

holm; P. Hoeflich, University of Texas; D. A. Howell, Lawrence Berkeley National Laboratory (LBNL); P. Lundqvist, Stockholm; L. Wang, LBNL; and J. C. Wheeler, University of Texas, report that SN 2001dm (cf. IAUC 7679) was observed with the Very Large Telescope UT3 (+ FORS1) at the ESO on Aug. 13 UT. A preliminary reduction of the spectrum (range 416–860 nm; spectral resolution 1.27 nm and 0.265 nm/pixel) reveals SN 2001dm to be a type-Ia supernova, ∼ 4–5 weeks past maximum brightness.

2001 August 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green